OpenStudy (anonymous):

Lance baked a pie. Justin ate 1/3 of the pie, and Joey ate 1/4 of the pie. How much of the pie is left?

OpenStudy (shamim):

1/3+1/4=?

OpenStudy (shamim):

Response plz!!

OpenStudy (anonymous):

I apologize; tech problems! How do I find the common denominator for the fractions? @shamim

OpenStudy (shamim):

Sorry..... i m also late

OpenStudy (shamim):

We can multiply by 12 both denominator n numerator

OpenStudy (shamim):

4/12+3/12=7/12 Right?

OpenStudy (anonymous):

Yes, thank you!

OpenStudy (shamim):

Now 1-7/12=?

OpenStudy (shamim):

U r most welcome!!!!

OpenStudy (shamim):

Medal???!!!!!